Ayala Inc. has conducted the following analysis related to its product lines, using a traditional costing system (volume-based) and an activity-based costing system. Both the traditional and the activity-based costing systems include direct materials and direct labor costs.

Total Costs

Products

Sales Revenue

Traditional

ABC

Product 540X $206,000 $57,000 $49,200
Product 137Y 156,000 49,000 27,000
Product 249S 93,000 18,000 47,800
For each product line, compute operating income using the traditional costing system.
Product 540X $

Product 137Y $

Product 249S $

LINK TO TEXT

For each product line, compute operating income using the activity-based costing system.
Product 540X $

Product 137Y $

Product 249S $

LINK TO TEXT

Using the following formula, compute the percentage difference in operating income for each of the product lines of Ayala: [Operating Income (ABC) – Operating Income (traditional cost)] ÷ Operating Income (traditional cost). (Round answers to 2 decimal places, e.g. 12.25. Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)
Product 540X

%
Product 137Y

%
Product 249S

%
Answer #1

Solution:

Operating income using the traditional costing system = Sales revenue - tradition costs

Product 540X = 206000 - 57000 = $149,000

Product 137Y = 156000 - 49000 = $107,000

Product 249S = 93000 - 18000 = $75,000

Operating income using the activity-based costing system = Sales revenue - ABC costs

Product 540X = 206000 - 49200 = $156,800

Product 137Y = 156000 - 27000 = $129,000

Product 249S = 93000 - 47800 = $45,200

Percentage difference in operating income = [Operating Income (ABC) – Operating Income (traditional cost)] ÷ Operating Income (traditional cost)

Product 540X = (156800- 149000) / 149000 = 5.23%

Product 137Y = (129000-107000) / 107000 = 20.56%

Product 249S = (45200 - 75000)/ 75000 = - 39.73%